The diameters of the main rotor and tail rotor of a single-engine helicopter are 7.53 m and 1.00 m, respectively. The respective rotational speeds are 451 rev/min and 4,150 rev/min. Calculate the speeds of the tips of both rotors.

main rotor ______m/s

tail rotor ______m/s

Compare these speeds with the speed of sound, 343 m/s.

vmain rotor = ______vsound

vtail rotor =________ vsound
Answer #1

speed = 451 rpm

angular speed of main rotor = 2pi*N / 60 = 2pi*451 / 60 = 47.22 rad/s

linear speed = radius * angular speed

linear speed of main rotor = (7.53/2) * 47.22 = 177.8 m/s

angular speed of tail rotor = 2pi*4150 / 60 = 434.6 rad/s

linear speed of tail rotor = (1/2) * 434.6 = 217.3 m/s

V(main rotor) / V(sound) = 177.8 / 343

V(main rotor) = 0.518 V(sound)

V(tail rotor) / V(sound) = 217.3 / 343

V(tail rotor) = 0.633 V(sound)